If the length of a rectangle is 1 foot less than twice its width and its perimeter is 34 feet, what is the length of the rectangle in feet?
Answer by Timnewman(323)

let the width of the rectangle be w
now,length of the
rectangle(L) =2w-1 -----(1)
perimeter of the rectangle
2(w+L)=34
divide both side by 2,you get
w+L=17 -----(2)
from (2)
w=17-w -----(3)
subsitute for w in (1)
that is, L=2(17-L)-1
clear the bracket
L=34-2L-1
Then, L+2L=33
3L=33
divide bothside by 3
3L/3=33/3
L=33/3
L=11ft
The lenght of the rectangle is 11ft.
HOPE THIS HELP,..